# Language Models Optimized to Fool Detectors Still Have a Distinct Style (And How to Change It)

### v1

@@ Abstract
Despite considerable progress in the development of machine-text detectors, it has been suggested that the problem is inherently hard, and therefore, that stakeholders should proceed under the assumption that machine-generated text cannot be reliably detected as such. We examine a recent such claim by Nicks et al. (2024) regarding the ease with which language models can be optimized to degrade the performance of machine-text detectors, including detectors not specifically optimized against. We identify a feature space—the stylistic feature space—that is robust to such optimization, and show that it may be used to reliably detect samples from language models optimized to prevent detection. Furthermore, we show that even when models are explicitly optimized against stylistic detectors, detection performance remains surprisingly unaffected. We then seek to understand if stylistic detectors are inherently more robust. To study this question, we explore a new paraphrasing approach that simultaneously aims to close the gap between human writing and machine writing in stylistic feature space while avoiding detection using traditional features. We show that when only a single sample is available for detection, this attack is universally effective across all detectors considered, including those that use writing style. However, as the number of samples available for detection grows, the human and machine distributions become distinguishable. This observation encourages us to introduce AURA, a metric that estimates the overlap between human and machine-generated distributions by analyzing how detector performance improves as more samples become available. Overall, our findings underscore previous recommendations to avoid reliance on machine-text detection. [^1] The datasets, method implementations, model checkpoints, and experimental scripts, will be released along with the paper: https://github.com/rrivera1849/style-aware-paraphrasing
@@ 1 Introduction
Large language models (LLMs) can generate fluent text across various domains. While there are many benign uses of LLMs, such as for writing assistance, they may also be abused (Weidinger et al., 2022; Hazell, 2023). To mitigate potential abuse, several machine-text detection systems have been proposed, including zero-shot methods such as Binoculars, DetectGPT, FastDetectGPT, and DNA-GPT (Hans et al., 2024; Mitchell et al., 2023; Bao et al., 2024; Yang et al., 2023), supervised detectors such as RADAR and OpenAI’s detector (Hu et al., 2023; Solaiman et al., 2019), and watermarking approaches (Kirchenbauer et al., 2024; Kuditipudi et al., 2024). However, as the gap between machine-generated and human-written text distributions narrows, detecting AI-generated text becomes increasingly challenging, raising concerns about the reliability of existing detection methods. Moreover, if this gap closes beyond a certain threshold, machine-text detection with acceptable false-positive rates may become difficult.
A recent study (Nicks et al., 2024) has shown that LLMs can be easily optimized to evade machine-text detectors by using a detector’s “humanness" score as a reward signal in reinforcement learning. However, while this approach defeats detectors that use token-level features of the predicted conditional distributions (Ippolito et al., 2020; Mitchell et al., 2023; Bao et al., 2024; Hans et al., 2024), we show that detectors that use writing style (Soto et al., 2024) remain robust to the distribution shift introduced during optimization. This suggests that the features used by these detectors are distinct from those indicative of writing style (Figure 1). Moreover, we find that style-based detectors remain robust even when targeted by optimization, an effect we attribute to the diversity of human writing styles. To robustly avoid detection and close the distributional gap, we argue that one must optimize both against detectors and for author-specific human writing styles—eliminating telltale signs easily spotted by detectors while also closing the gap between human and machine text writing styles.
@@
StyleDPOOurs (§missing $\mathcal{M}$$\mathcal{H}$Nicks et al. (2024)Ours (§missing 3)
@@
Figure 1: (a) UMAP (McInnes et al., 2020) projections of representations that capture writing style for comments in the Reddit, using LUAR (Rivera-Soto et al., 2021). Each point corresponds to a document of at most 128 tokens. Despite optimization against FastDetectGPT, the LLM’s writing style remains largely unchanged (compare with ). In contrast, our approach better closes the gap between human-written and machine-generated text (compare with ). (b) Cartoon version of (a) illustrating our main findings where $\mathcal{M}$ denotes the distribution of machine-generated text and $\mathcal{H}$ the distribution of human-written text. Here, we illustrate that stylistic space separates DPO-optimized LLM samples from human text (§missing 2); and that stylistic-paraphrasing closes the gap between human and machine-generated text (§missing 3).
To better understand the limits of detectability, we measure the area under the receiver operating curve (AUROC) as the sample size increases, finding that it increases as more samples are provided—supporting the findings of (Chakraborty et al., 2023) linking increases in AUROC to the sample size and the total variation (TV). Building on the theoretical link established by Chakraborty et al. (2023), we introduce AURA, a metric that tracks how the estimated maximum distributional overlap evolves with increasing samples, and provides the area under this curve as a single scalar summary of the similarity between the distributions as observed by a machine-text detector. We find that while approaches that rely on preference optimization help evade detection, the gap between the distributions remains large, implying that most modern detectors rely on surface level features that are easy to alter.
@@
We show that although LLMs can be optimized to defeat machine-text detectors, they remain identifiable by their writing style (§missing 2). We introduce a training recipe for a state-of-the-art style-aware paraphraser that mimics human writing style while evading machine-text detectors (§missing 3). We extend the work of Chakraborty et al. (2023) by proposing a sample-based measure, AURA, which quantifies the upper bound of the overlap between human- and machine-text distributions (§missing 4).
@@ 2 Stylistic Detectors are Robust Against Optimization
Table 1: Machine-text detection performance (AUROC) of various detectors evaluated on outputs from Mistral-7B, with and without optimization against machine-text detectors. While optimization against FastDetectGPT (Mistral-7B-DPO-FastDetectGPT) significantly degrades the performance of both FastDetectGPT and Binoculars, StyleDetect remains robust. Optimizing against StyleDetect (Mistral-7B-DPO-StyleDetect) does not reduce its performance, suggesting that DPO is insufficient to close the gap between the writing styles. Experiments on more LLMs are reported in §missing 6.
@@
In this section, we show that machine-text detectors that use features indicative of writing style are robust against optimization. Recently, Nicks et al. (2024) showed that LLMs can be easily optimized to evade machine-text detectors by using a detector’s “humanness” score as a reward signal in reinforcement learning. This strategy was shown to significantly degrade the performance of detectors that rely on features derived from the predicted conditional distributions such as FastDetectGPT (Bao et al., 2024) and Binoculars  (Hans et al., 2024). However, it remains unclear whether detectors that use writing style, such as that proposed by Soto et al. (2024), exhibit the same vulnerability to optimization. To test the robustness of such detectors, we optimize Mistral-7B to generate responses to Reddit comments that are rated as more human-like by FastDetectGPT. We also perform optimization against the style-based detector proposed by Soto et al. (2024), which we refer to as StyleDetect. Since StyleDetect requires exemplars from the machine class, we provide $1$ example from the unoptimized Mistral-7B model. [^2] More than one exemplar can be used to improve performance; performance saturates at $5$ examples in this experiment. Its detection score is the cosine similarity between a test sample and the single machine-example in the stylistic embedding space. We evaluate each detector using the AUROC, results are shown in Table 1. When we optimize Mistral-7B against FastDetectGPT, the AUROC for both Binoculars and FastDetectGPT drops below random. In contrast, we observe that StyleDetect remains robust, with no drop in AUROC, which implies that after optimization the writing style of Mistral-7B remains largely unchanged. When optimizing against StyleDetect, we observe no degradation in the performance of StyleDetect, but observe degradation of both Binoculars and FastDetectGPT. These results suggest that the features indicative of writing style are distinct from those used by detectors that use features derived from the predicted conditional distributions. We attribute the difficulty of optimizing against StyleDetect to the variability found in human writing styles. For example, consider one author who writes exclusively in capital letters and another who writes entirely in lowercase. Simultaneously optimizing for both styles creates conflicting objectives, making it difficult for the model to converge to a good solution.
@@ 3 Building a Hard to Detect Style-Aware Paraphraser
Mistral-7B-DPO-FastDetectGPT Wishing for ’Ramsay Snow’ mention or Jon’s retort, yet appreciating the dogs’ presence instead. A testament to George R.R. Martin’s captivating narrative.
@@
Given a machine-generated text sample, our goal is to produce a paraphrase that closely mimics the writing style of a human author. However, parallel data that maps machine-generated text to its human-written paraphrase does not exist. Hence, we first build a paraphraser that, given $M$ in-context pairs of machine-generated paraphrases and their human-written originals, maps a new paraphrase back to its original. Such data can be readily generated, for example, by paraphrasing human-written text with an LLM. Formally, given a dataset of human-written texts $x_{i}$, their machine-generated paraphrases $p_{i}$, and their corresponding author labels $a_{i}$, denoted as $\mathcal{D}_{para}=\{(x_{i},p_{i},a_{i})\}_{i=1}^{N}$, we instruction-tune (Wei et al., 2022) [^3] Instruction can be found in §missing F.4 an LLM to model $p(x_{i}\mid p_{i},C_{i})$ where $C_{i}=\{(x_{j},p_{j}):a_{j}=a_{i},j\neq i\}$ are exemplars pairs (original and paraphrases) from the same author. In practice, for each human-written text $x_{i}$ we generate $P$ paraphrases, adding all $P*M$ exemplars to the context. Generating multiple paraphrases per human-written text is an efficient way to increase the number of exemplars without incurring the additional cost of collecting more human-written samples.
@@
4 AURA: Measuring the Distributional Gap
In our experiments, we find that when detectors are provided with a single sample, our systems generations are indistinguishable from human-written text. However, Chakraborty et al. (2023) find that the distributions become more distinct as more samples are provided, establishing a theoretical lower bound on the number of samples $n$ required for a detector to achieve AUROC $\epsilon$, given the total variation $\delta$ between human- and machine-text distributions:
$n\geq\frac{1}{\delta^{2}}\log\left(\frac{2}{1-\epsilon}\right)$ (1)
This result encourages us to estimate the maximum possible overlap between the two distributions as a function of $n$ and the AUROC achieved when $n$ samples are provided $\epsilon_{n}$, we denote this function as $S(n,\epsilon_{n})$. Solving (missing) 1 for $1-\delta$ gives:
$S(n,\epsilon_{n})=1-\delta\leq 1-\sqrt{\frac{\log\left(\frac{2}{1-\epsilon_{n}}\right)}{n}}$ (2)
By computing $S(n,\epsilon_{n})$ for increasing $n$, we obtain a curve $\mathcal{C}=\{(n,S(n,\epsilon_{n}))\}_{n=1}^{N}$, where $N$ is the total number of samples available. Our proposed metric, $\texttt{AURA}=\int S(n,\epsilon_{n})$, is the area under this curve (AUC). It is a quantitative measure of the maximum distributional overlap, with higher values indicating greater similarity between human and machine distributions as assessed by the detector. [^4] The proposed measure is similar to measures used in kernel two-sample tests, such as maximum mean discrepancy Gretton et al. (2012), but is tailored to our machine-text detection task. To derive a single score from multiple samples, we aggregate the detector’s outputs by computing the sample mean across $n$ samples. This score is then used to compute the AUROC.
@@ 5.1 Datasets
For methods that require preference data, namely ours and Mistral-7B-DPO-FastDetectGPT, we subsample additional text from each domain, including Reddit, Amazon reviews (Ni et al., 2019), and Blogs (Schler et al., 2006). Specifically, we draw $10{,}000$ samples each from unique authors in the Reddit and Amazon datasets, and $6{,}000$ from the Blogs dataset, ensuring all authors are distinct and disjoint from those in training and evaluation sets.
@@
We evaluate our approach across three domains: Reddit, Amazon reviews, and Blogs. To generate machine text, we prompt one of Mistral-7B, gpt-4o-mini, or Llama-3, chosen uniformly at random, to create new comments, reviews, or blog snippets (see prompts in Appendix F). Each baseline described in §missing 5.2 is then applied to modify this generated text to evade detection. The only exception is Mistral-7B-DPO-FastDetectGPT, which generates the text directly, rather than modifying pre-existing outputs. For methods that require target exemplars, including our own, we randomly select an author from the dataset to define the target style and provide $16$ of their texts as exemplars.
@@ 5.2 Baselines
We prompt gpt-4o-mini to paraphrase machine-generated text. Paraphrasing has been shown to be an effective attack against detectors (Krishna et al., 2023; Sadasivan et al., 2025; Soto et al., 2025), as it alters surface-level features while preserving semantic contents.
@@
is a lightweight (800M parameter) style-aware paraphraser trained on Reddit that uses pre-trained author representations for efficient few-shot style transfer (Horvitz et al., 2024b). In contrast, our system tunes a Mistral-7B with LoRA (Rafailov et al., 2024), does not rely on author representations, and is explicitly optimized to evade machine-text detectors.
@@ 5.3 Metrics, Detectors, and Inference
To measure the performance of machine-text detectors, we use the standard area under the curve of the receiver operating curve, referred to as AUROC. To better align with real-world scenarios where false-positives are costly, we calculate the partial area for FPRs less than or equal to 10%, which we refer to as AUROC(10). To measure how well the meaning of text is preserved after modification, we use SBERT555sentence-transformers/all-mpnet-base-v2, computing the cosine similarity between embeddings of the original and modified text. Finally, to measure how well the style-aware paraphrasing methods introduce the target style, we use CISR [^6] AnnaWegmann/Style-Embedding, computing the cosine similarity between embeddings of the generated text and target exemplars.
@@
To evaluate how detectable our generations are, we use various detectors, including Rank (Gehrmann et al., 2019), LogRank (Solaiman et al., 2019), FastDetectGPT (Bao et al., 2024), Binoculars (Hans et al., 2024), OpenAI’s detector (Solaiman et al., 2019), RADAR (Hu et al., 2023), and StyleDetect (Soto et al., 2024). For FastDetectGPT, we use gpt-neo-2.7B, the backbone originally used by the authors. For Rank and LogRank, we use gpt2-xl as the backbone. StyleDetect operates in a few-shot setting, requiring exemplars from the machine-text class; we provide $K=100$ such examples drawn from random machine-generated text in our dataset that was not produced by any of the evaluated methods.
@@
During inference, our goal is to paraphrase a fully machine-generated sample in the style of a human-author. However, during training, only machine paraphrases of human text were observed. This introduces a distribution mismatch, as our system was trained on paraphrases of human-text, which oftentimes contain tokens copied from the original human-text. To bridge this gap, we iteratively apply our style-aware paraphraser, gradually reducing the distributional mismatch. At each iteration, we generate $10$ candidates, and choose the top-$P$ that best preserves semantics according to SBERT777sentence-transformers/all-mpnet-base-v2 for the next iteration. In the final iteration, we simply pick the best candidate according to meaning preservation. When our system is applied to paraphrases of human-written text, we simply generate one candidate generation.

### v2

@@ Abstract
Despite considerable progress in the development of machine-text detectors, it has been suggested that the problem is inherently hard, and therefore, that stakeholders should proceed under the assumption that machine-generated text cannot be reliably detected as such. We examine a recent such claim by Nicks et al. (2024) regarding the ease with which language models can be optimized to degrade the performance of machine-text detectors, including detectors not specifically optimized against. We identify a feature space—the stylistic feature space—that is robust to such optimization, and show that it may be used to reliably detect samples from language models explicitly optimized to prevent detection. Furthermore, we show that even when models are explicitly optimized against stylistic detectors, detection performance remains surprisingly unaffected. We then seek to understand if stylistic detectors are inherently more robust. To study this question, we explore a new paraphrasing approach that simultaneously aims to close the gap between human writing and machine writing in stylistic feature space while avoiding detection using traditional features. We show that when only a single sample is available for detection, this attack is universally effective across all detectors considered, including those that use writing style. However, as the number of samples available for detection grows, the human and machine distributions become distinguishable. Overall, our findings underscore previous recommendations to avoid reliance on machine-text detection on individual documents. [^1] The datasets, method implementations, model checkpoints, and experimental scripts, will be released along with the paper: https://anonymous.4open.science/status/style-aware-paraphrasing-BD8E
@@ 1 Introduction
Large language models (LLMs) can generate fluent text across various domains. While there are many benign uses of LLMs, such as for writing assistance, they may also be abused (Weidinger et al., 2022; Hazell, 2023). To mitigate potential abuse, several machine-text detection systems have been proposed, including zero-shot methods such as Binoculars, DetectGPT, FastDetectGPT, and DNA-GPT (Hans et al., 2024; Mitchell et al., 2023; Bao et al., 2024; Yang et al., 2023), supervised detectors such as RADAR and ReMoDetect (Hu et al., 2023; Lee et al., 2024), and watermarking approaches (Kirchenbauer et al., 2024; Kuditipudi et al., 2024). However, as the gap between machine-generated and human-written text distributions narrows, detecting AI-generated text becomes increasingly challenging, raising concerns about the reliability of existing detection methods. Moreover, if this gap closes beyond a certain threshold, machine-text detection with acceptable false-positive rates may become difficult.
Recently, Nicks et al. (2024) has shown that LLMs can be easily optimized to evade machine-text detectors by using a detector’s “humanness" score as a reward signal in reinforcement learning. However, while this approach defeats detectors that use token-level features of the predicted conditional distributions (Ippolito et al., 2020; Mitchell et al., 2023; Bao et al., 2024; Hans et al., 2024), we show that detectors that use writing style (Soto et al., 2024) remain robust to the distribution shift introduced during optimization. This suggests that the features used by these detectors are distinct from those indicative of writing style (Figure 1). Moreover, we find that style-based detectors remain robust even when targeted by optimization, an effect we attribute to the diversity of human writing styles. To robustly avoid detection and close the distributional gap, we argue that one must optimize both against detectors and for author-specific human writing styles—eliminating telltale signs easily spotted by detectors while also closing the gap between human and machine text writing styles.
@@
[…]
@@
Figure 1: (a) UMAP (McInnes et al., 2020) projections of representations that capture writing style for comments in the Reddit domain, using LUAR (Rivera-Soto et al., 2021). Each point corresponds to a document of at most 128 tokens. Despite optimization against FastDetectGPT, the LLM’s writing style remains largely unchanged (compare with ). In contrast, our approach better closes the gap between human-written and machine-generated text (compare with ). (b) Cartoon version of (a) illustrating our main findings where $\mathcal{M}$ denotes the distribution of machine-generated text and $\mathcal{H}$ the distribution of human-written text. Here, we illustrate that stylistic space separates DPO-optimized LLM samples from human text (§ 2); and that stylistic-paraphrasing closes the gap between human and machine-generated text (§ 3).
To better understand the limits of detectability, we measure the area under the receiver operating curve (AUROC) as the sample size increases, finding that it increases as more samples are provided—supporting the findings of (Chakraborty et al., 2023) linking increases in AUROC to the sample size and the total variation (TV). We find that while approaches that rely on preference optimization help evade detection, the gap between the distributions remains large, implying that most modern detectors rely on surface level features that are easy to alter.
@@
We show that although LLMs can be optimized to defeat machine-text detectors, they remain identifiable by detectors that avail of writing style and that moreover, optimizing against such detectors does not reduce their performance.(§ 2). We introduce a training recipe for a state-of-the-art style-aware paraphraser that mimics human writing style while evading machine-text detectors (§ 3).
@@ 2 Stylistic Detectors are Robust Against Optimization
Qwen-7B-Instruct 0.47 0.50 0.98
Qwen-7B-Instruct-DPO-FastDetectGPT 0.49 0.53 0.97
Qwen-7B-Instruct-DPO-StyleDetect 0.47 0.54 0.97
Mistral-Nemo-Instruct 0.75 0.79 0.97
Mistral-Nemo-Instruct-DPO-FastDetectGPT 0.37 0.33 0.96
Mistral-Nemo-Instruct-DPO-StyleDetect 0.67 0.67 0.95
@@
Table 1: Machine-text detection performance (AUROC) of various detectors evaluated on outputs from Mistral-7B, Qwen-7B, and Mistral-Nemo with and without optimization against machine-text detectors. While optimization against FastDetectGPT (variants with -DPO-FastDetectGPT suffix) significantly degrades the performance of both FastDetectGPT and Binoculars, StyleDetect remains robust. Optimizing against StyleDetect (variatns with -DPO-StyleDetect suffix) does not reduce its performance, suggesting that DPO is insufficient to close the gap between the writing styles. Experiments on more LLMs are reported in § 5.
@@
In this section, we show that machine-text detectors that use features indicative of writing style are robust against optimization. Recently, Nicks et al. (2024) showed that LLMs can be easily optimized to evade machine-text detectors by using a detector’s “humanness” score as a reward signal in reinforcement learning. This strategy was shown to significantly degrade the performance of detectors that rely on features derived from the predicted conditional distributions such as FastDetectGPT (Bao et al., 2024) and Binoculars  (Hans et al., 2024). However, it remains unclear whether detectors that use writing style, such as that proposed by Soto et al. (2024), exhibit the same vulnerability to optimization. To test the robustness of such detectors, we optimize Mistral-7B, Qwen-7B, and Mistral-Nemo to generate responses to Reddit comments that are rated as more human-like by FastDetectGPT. We also perform optimization against the writing-style-based detector proposed by Soto et al. (2024), which we refer to as StyleDetect. Since StyleDetect requires exemplars from the machine class, we provide $100$ examples from the unoptimized LLM model. Its detection score is the cosine similarity between a test sample and the averaged embedding of the $100$ machine-examples in the stylistic embedding space. We evaluate each detector using the AUROC showing results in Table 1. When we optimize any of the LLMs against FastDetectGPT, the AUROC of both Binoculars and FastDetectGPT drops below random in cases where FastDetectGPT was originally discriminative (Mistral-7B and Mistral-Nemo). In contrast, we observe that StyleDetect remains robust, with no significant drop in AUROC, which implies that after optimization the writing style of each LLM remains largely unchanged. When optimizing against StyleDetect, we observe no significant degradation in the performance of StyleDetect. These results suggest that the features indicative of writing style are distinct from those used by detectors that use features derived from the predicted conditional distributions. We attribute the difficulty of optimizing against StyleDetect to the variability found in human writing styles. For example, consider one author who writes exclusively in capital letters and another who writes entirely in lowercase. Simultaneously optimizing for both styles creates conflicting objectives, making it difficult for the model to converge to a good solution.
@@ 3 Building a Hard to Detect Style-Aware Paraphraser
Mistral-7B-DPO-FastDetectGPT Wishing for ‘Ramsay Snow’ mention or Jon’s retort, yet appreciating the dogs’ presence instead. A testament to George R.R. Martin’s captivating narrative.
@@
Given a machine-generated text sample, our goal is to produce a paraphrase that closely mimics the writing style of a human author. However, parallel data that maps machine-generated text to its human-written paraphrase does not exist. Hence, we first build a paraphraser that, given $M$ in-context pairs of machine-generated paraphrases and their human-written originals, maps a new paraphrase back to its original. Such data can be readily generated, for example, by paraphrasing human-written text with an LLM. Formally, given a dataset of human-written texts $x_{i}$, their machine-generated paraphrases $p_{i}$, and their corresponding author labels $a_{i}$, denoted as $\mathcal{D}_{para}=\{(x_{i},p_{i},a_{i})\}_{i=1}^{N}$, we instruction-tune (Wei et al., 2022) [^2] Instruction can be found in § F.4 an LLM to model $p(x_{i}\mid p_{i},C_{i})$ where $C_{i}=\{(x_{j},p_{j}):a_{j}=a_{i},j\neq i\}$ are exemplars pairs (original and paraphrases) from the same author. In practice, for each human-written text $x_{i}$ we generate $P$ paraphrases, adding all PMP*M exemplars to the context. Generating multiple paraphrases per human-written text is an efficient way to increase the number of exemplars without incurring the additional cost of collecting more human-written samples.
@@ 4.1 Datasets
For methods that require preference data, namely ours and Mistral-7B-DPO-FastDetectGPT, we subsample additional text from each domain, including Reddit, Amazon reviews (Ni et al., 2019), and Blogs (Schler et al., 2006). Specifically, we draw 10,00010{,}000 samples each from unique authors in the Reddit and Amazon datasets, and 6,0006{,}000 from the Blogs dataset, ensuring all authors are distinct and disjoint from those in training and evaluation sets.
@@
We evaluate our approach across three domains: Reddit, Amazon reviews, and Blogs. To generate machine text, we prompt one of Mistral-7B, gpt-4o-mini, or Llama-3, chosen uniformly at random, to create new comments, reviews, or blog snippets (see prompts in Appendix F). Each baseline described in § 4.2 is then applied to modify this generated text to evade detection. The only exception is Mistral-7B-DPO-FastDetectGPT, which generates the text directly, rather than modifying pre-existing outputs. For methods that require target exemplars, including our own, we randomly select an author from the dataset to define the target style and provide $16$ of their texts as exemplars.
@@ 4.2 Baselines
Paraphrasing has been shown to be an effective attack against detectors (Krishna et al., 2023; Sadasivan et al., 2025; Soto et al., 2025), as it alters surface-level features while preserving semantic contents. As such, we evaluate against two paraphrasing baselines. Our first paraphrasing baseline prompts gpt-4o-mini to paraphrase machine-generated text. Our second baseline uses DIPPER (Krishna et al., 2023), an 11 billion parameter paraphrasing model built to evade detectors.
OUTFOX
is an attack that incorporates in-context examples of text detected as human or machine by a detector, prompting the LLM to generate text that would be detected as human (Koike et al., 2024). We chose to include $16$ text samples along with the detection results of StyleDetect (instantiated with $100$ few-shot samples). This attack is significant in that it evaluates whether or not prompting is enough to close the gap between human-written and machine-generated styles.
@@
is a lightweight (800M parameter) style-aware paraphraser trained on Reddit that uses pre-trained author representations for efficient few-shot style transfer (Horvitz et al., 2024b). In contrast, our system tunes a Mistral-7B with LoRA (Hu et al., 2021), does not rely on author representations, and is explicitly optimized to evade machine-text detectors.
@@ 4.3 Metrics, Detectors, and Inference
To measure the performance of machine-text detectors, we use the standard area under the curve of the receiver operating curve, referred to as AUROC. To better align with real-world scenarios where false-positives are costly, we calculate the partial area for FPRs less than or equal to 10%, which we refer to as AUROC(10). To measure how well the meaning of text is preserved after modification, we use SBERT333sentence-transformers/all-mpnet-base-v2, computing the cosine similarity between embeddings of the original and modified text. Finally, to measure how well the style-aware paraphrasing methods introduce the target style, we use CISR [^4] AnnaWegmann/Style-Embedding, computing the cosine similarity between embeddings of the generated text and target exemplars.
@@
To evaluate how detectable our generations are, we use various detectors, including Rank (Gehrmann et al., 2019), LogRank (Solaiman et al., 2019), FastDetectGPT (Bao et al., 2024), Binoculars (Hans et al., 2024), ReMoDetect (Lee et al., 2024), RADAR (Hu et al., 2023), and StyleDetect (Soto et al., 2024). For FastDetectGPT, we use gpt-neo-2.7B, the backbone originally used by the authors. For Rank and LogRank, we use gpt2-xl as the backbone. StyleDetect operates in a few-shot setting, requiring exemplars from the machine-text class; we provide $K=100$ such examples drawn from random machine-generated text in our dataset that was not produced by any of the evaluated methods. Moreover, we include two additional versions of StyleDetect that rely on different underlying stylistic representations. We also include two StyleDetect variants that use different style representations: one with CISR [^5] AnnaWegmann/Style-Embedding embeddings (StyleDetect-CISR) and another with StyleDistance [^6] StyleDistance/styledistance embeddings (StyleDetect-SD). In total, we evaluate nine detectors across trained classifiers (RADAR, ReMoDetect), zero-shot detectors (Rank, LogRank, FastDetectGPT, Binoculars), and few-shot stylistic detectors (StyleDetect, StyleDetect-CISR, StyleDetect-SD).
@@
To defeat detection, our goal is to paraphrase a fully machine-generated sample in the style of a human-author. However, during training, only machine paraphrases of human text were observed. This introduces a distribution mismatch, as our system was trained on paraphrases of human-text, which oftentimes contain tokens copied from the original human-text. To bridge this gap, we iteratively apply our style-aware paraphraser, gradually reducing the distributional mismatch. At each iteration, we generate $10$ candidates, and choose the top-$P$ (number of paraphrases ingested by our system) that best preserve the semantics of the original text according to SBERT777sentence-transformers/all-mpnet-base-v2 for the next iteration. In the final iteration, we simply pick the candidate that best preserves the meaning of the original text. When our system is applied to paraphrases of human-written text, we simply generate one candidate generation.